Status Report

NASA ISS On-Orbit Status 31 January 2009

By SpaceRef Editor
February 1, 2009
Filed under , , ,
NASA ISS On-Orbit Status 31 January 2009

All ISS systems continue to function nominally, except those noted previously or below. Saturday – rest day for CDR Fincke, FE-1 Lonchakov & FE-2 Magnus.

Fincke & Magnus started the day with their daily download of the accumulated data of the SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment from their Actiwatches to the HRF-1 (Human Research Facility 1) laptop as part of another week-long session with SLEEP, their second. [To monitor the crewmember’s sleep/wake patterns and light exposure, the crewmembers wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew’s discretionary “job jar” task list. It is the third session for Mike, the second for Sandra.]

Lonchakov performed his part of the regular weekly three-hour task of thorough station cleaning, focusing on the RS (Russian Segment). ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

In the SM, the FE-1 later also completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers, replacement of EDV-SV waste water and EDV-U urine containers and performing US condensate processing (transfer from CWC to EDV containers) if condensate is available.]

For Mike & Sandy, most of their day’s schedule was taken up by performing the first BRASLET SDTO (Station Development Test Objective) protocol, with both of them subjects and assisting each other in turn during the scan sessions. [Prior to the sessions, the protocol had called for no BRASLET cuff within 24 hrs, no caffeine within 12 hrs, no heavy meals within 4 hrs, no food or exercise at all within 2 hrs, and no liquids within 30 min. Background: BRASLET (Validation of On-Orbit Methodology for the Assessment of Cardiac Function and Changes in the Circulating Volume Using Ultrasound and BRASLET-M Occlusion Cuffs) is SDTO 17011, sponsored by NASA and FSA/IBMP (Russian Federal Space Agency/Institute of Bio-Medical Problems, Russian: IMBP, Dr. Valery Bogomolov). BRASLET is testing the performance of occlusion cuffs in modifying fluid shifts that occur early during physiological transition into the space environment. Understanding the effects of this countermeasure on cardiovascular function will be useful for both medical operations and future research. The goal of this investigation is to establish a valid ultrasound methodology for assessing a number of aspects of central and peripheral hemodynamics and cardiovascular function, specifically in rapid changes in intravascular circulating volume. The SDTO uses BRASLET-M occlusion cuffs, which are a Russian-made operational countermeasure already pre-calibrated and available onboard for each ISS crewmember. BRASLET uses multiple modes of ultrasound imaging and measurements, in combination with short-term application of BRASLET-M occlusive cuffs and cardiopulmonary maneuvers (Valsalva, Mueller) to demonstrate and to evaluate the degree of changes in the circulating volume on orbit. This will be accomplished by performing echocardiographic examinations in multiple modes (including Tissue Doppler mode), ultrasound measurements of lower extremity venous and arterial vascular responses to BRASLET-M device under nominal conditions and also during cardiopulmonary Mueller and Valsalva maneuvers. Identical measurements are being repeated without BRASLET-M, with BRASLET-M applied, and immediately after releasing the occlusion device.]

At ~9:00am EST, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

Sandy completed the daily flushing of the ER6 PWD (EXPRESS Rack 6/Potable Water Dispenser), now from a CWC-I (Contingency Water Container-Iodinated, #1013, relabeled as Special Fluids) instead of drink bags. [The PWD had been found, via several microbial analyses by Magnus, to have bacteria growing in the ambient water leg. Latest microbial results indicate that not enough iodine may get into the system to kill off any microbes, since the amount of 250 mL used lately did not take into account the filter and, as ground testing has shown, it takes about 24 hours for the iodine to convert to non-biocidal iodide when left stagnant in a filter like the one used in the PWD. The amount of iodinated flush water was reduced last Friday (1/30) to 0.5 L for the next seven days (to end 2/6).]

At ~12:55pm EST, Lonchakov powered up the SM’s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply), including the RSK2 laptop with the WinPack 6.8 ham radio application, and at 1:15pm conducted with Fincke a radio exchange with school kids at Montreal, Canada.

The station residents completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), ARED advanced resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Two voluntary tasks have been added to Mike & Sandy’s “job jar” task list for their choice: (a) an EVA bag audit & reconfiguration for the 15A spacewalks; (b) more troubleshooting on the G1 camcorder MPC HD video system which is creating horizontal “artifacts” during downlink. [The troubleshooting should identify if it is a problem with the HD camcorder or the tape that was used.]

TOCA Update: The TOCA (Total Organic Carbon Analyzer) run to analyze a WPA (Water Processor Assembly) water sample reported yesterday was aborted due to a low pressure indication (“P1 < 4.0 psig and in line flush"). Engineering team is investigating.

WHC Update: The recent (1/29) failure indication of the Dose Pump LED (light-emitting diode) of the US/Russian WHC (Waste & Hygiene Compartment) has been narrowed down by NASA & RSC-Energia specialists to a position sensor issue in the Dose Pump or a logic corruption in the ASU CP (Control Panel). Engineers do not want to replace either the pump or the ASU CP until the failure is better understood, to avoid the risk of wasting spare hardware. A troubleshooting activity with the crew is scheduled on 2/3.

Weekly Science Update (Expedition Eighteen — Week 14)

3-D SPACE: Complete 3 in 1J Stage. Planned: 4th in Inc-19.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): Planned/Reserve.

BIO-4: Complete.

BIOLAB (BLB): Activities to prepare the BIOLAB facility fully in view of the next experiment in Incs-19/20 continued with some ground activities for maintenance file uplink and the HM (Handling Mechanism) positioning and functional test which were executed nominally on 1/21.

CARDIOCOG-2: Complete.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Mike, from your request we are currently planning a combined session with you and Sandy as a Voluntary Science option for 2/7. This would be an optional mid-increment session. You would still be required to complete another session three to two weeks prior to landing to count as a subject. Sandy, we are currently planning your final on-orbit CCISS session to be offered through a Voluntary Science option on 2/7. This session has a constraint window of Return minus 21-14 days.”

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Ongoing.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

DomeGene (JAXA): Planned.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

EDR (European Drawer Rack): Planned. A smoke detector test was successfully carried out on 1/7.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): ENose is operating normally. Confirmational event done on 1/21 showed up well in ENose data.

EPM (European Physiology Module): MEEMM (Multi Electrodes Encephalogram Measurement Module) calibration was successfully performed on 1/21.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): Science data acquisition for the EuTEF instruments continues. EVC (Earth Viewing Camera) has been able to take pictures this week as the temperature has been high enough.

FSL (Fluid Science Laboratory): “Thanks for the performance of the troubleshooting activity on the FSL video tape recorder on 1/20-20. We hope a positive outcome on the noise can be confirmed when the FSL will be activated on 1/23.”

GEOFLOW: Aborted. As reported last week, a leak inside the concentric rotating spheres has caused investigators to abort this experiment. “Thank you for the performance on the successful troubleshooting on the FSL Video Tape Recorder. We are glad the beeping is no longer a bother.”

HDTV System Test DL (JAXA): Complete.

ICE CRYSTAL (JAXA): On-going. During the Ice Crystal experiment on 1/29, there was an anomaly. One of the temperature controls did not work as planned. It was assumed that the SCOF facility’s Electronic Controller had a temporal failure, so the SCOF was deactivated, and activated again. After re-activation, the temperature control was nominal, so the experiment is now nominal.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Reserve. “Thanks for operating LOCAD and being the first to use Gram+ cartridges aboard the ISS. These cartridges detect "Gram-positive" bacteria, a group of bacteria predominant on spacecraft cabin surfaces and which test ‘positive’ with the Gram stain (developed by Danish microbiologist Hans Christian Gram in 1884). The Gram+ LOCAD cartridges provide a miniaturized molecular test for Gram-positive bacteria. Your results on GMT024 showed that all four JEM PM surface sites contain low/zero levels of Gram-positive bacteria; similar to the low/zero levels found pre-flight at KSC. We shall share this data with our JAXA colleagues, who will be happy to learn that the JEM PM has remained ‘clean’ from a biological perspective. Thanks again for volunteering to operate LOCAD.”

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

Micro-G Clay (JAXA EPO): Complete.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): Complete.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY: “Mike thanks for the great work and attention to detail in completing your FD120 session within the timeline constraints! Your next session will be your FD180. Nice job working through the tacho error. Sandy thanks for sleep logging this week. We are planning another week of logging next week. Your only remaining activities would then be one more Actiwatch download and then doffing the Actiwatch on the Shuttle with the other crewmembers.”

PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

RadGene & LOH (JAXA): Planned. “On 1/26, MLT laptop terminal was configured to the measurement configuration. We had some trouble with the Ethernet connection, but with a reboot of the MLT, it is now nominal and ready to measure for the RadGene and LOH experiment. Thank you for your support.”

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): Complete. “Thank you, Mike, for all the hard work you have done on SHERE. Good job on your careful and meticulous packing of the SHERE hardware for future use – we very much appreciate it. A review of the HD video you provided of the test runs continues to provide further insight and information into the fluid’s behavior, and it has proved to be a wonderful addition to the data that was originally planned for SHERE. Best wishes to you on the rest of Increment 18!”

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Mike thanks for sleep logging this week. We will schedule your last required week later in the increment. Any additional logging is above and beyond and greatly appreciated by the PI.”

SOLAR (Solar Monitoring Observatory): The Sun observation window which started on 1/14 ended on 1/28. The measurements for SOLACES and SOLSPEC went well. The SOLAR platform went into Stand-by Mode from an anomaly on 1/26 where it was recovered to Pointing Mode. The anomaly occurred again on 1/28, but this time it was recovered to Idle Mode since the sun observation window had ended.

SOLO (Sodium Loading in Microgravity): Complete (1J Stage).

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VLE (Video Lessons ESA): “Thanks Sandy & Mike for the smooth performance of the VLE-1 Single meal activity on 1/9.”

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).

CEO (Crew Earth Observations): Ongoing.

CEO photo targets uplinked for today were East Haruj Megafans, Africa (weather was predicted be clear over these megafans, formed by rivers that once drained the Tibesti Mountains. Looking for sinuous stream channel patterns that overlap. Overlapping mapping frames, taken along track, were requested to build detailed maps of the megafans), Mississippi Delta Region, Louisiana (weather was predicted to be clear over this well-known delta. The current delta configuration is the result of human alteration of the Mississippi River through control structures, levees, and channelization. This has resulted in land loss along the Louisiana coastline that continues today. Overlapping frames, taken right of track, of the coastline were requested for comparison to historical imagery), Middlesboro Impact Crater, Kentucky (ISS had a near-nadir pass over this 6 km diameter impact structure. Development in the region has obscured much of the crater, but a rim can still be recognized. Overlapping frames, taken slightly to the right of track, were recommended to capture highly detailed imagery of the impact structure), and Madrean Sky Islands, southern AZ/northern Mexico (looking to the right of track for a southeast-looking view of the southern Basin and Range of Arizona, and the northwestern Sierra Madre of Mexico. The summits of these high mountain ranges include forested "sky islands" – remnants of extensive forests that once existed at lower elevations during cooler and wetter climate periods. Context imagery of the southern AZ and northwestern Mexico ranges was requested as an aid to locating particular sky island targets).

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 6:59am EST [= epoch]):
Mean altitude — 356.9 km
Apogee height — 362.2 km
Perigee height — 351.5 km
Period — 91.68 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0007954
Solar Beta Angle — -30.8 deg (magnitude increasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 29 m
Revolutions since FGB/Zarya launch (Nov. 98) — 58445

Significant Events Ahead (all dates Eastern Time, some changes possible!):
02/02/09 — IMMT Go/No-Go on 2/4 Reboost
02/04/09 — ISS Reboost (~2:50am EST), SM mid-ring thrusters, 0.5 m/s (not SM ME twin thrusters)
02/05/09 — Progress M-01M/31P undocking & deorbit (~11:00pm EST)
02/10/09 — Progress 32P launch
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment (7:32am EST)
02/13/09 — Progress 32P docking (2:20am EST); [crew wake: 10:30pm on 2/12]
02/14/09 — STS-119/Discovery/15A docking (3:57am EST)
02/23/09 — STS-119/Discovery/15A undocking (9:30pm EST)
02/26/09 — STS-119/Discovery/15A landing (KSC, 1:50am EST)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 — Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress 32P undocking & deorbit
05/12/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/27/09 — Soyuz TMA-15/19S launch
Six-person crew on ISS
08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC, last crew rotation
08/XX/09 — Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz
09/XX/09 — H-IIB (JAXA HTV-1)
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4
12/XX/11– Proton 3R/MLM w/ERA.

SpaceRef staff editor.